OpenStudy (anonymous):

The hypotenuse of an isosceles right triangle measures 15 feet. Find the area (?)

OpenStudy (anonymous):

a^2 + a^2 = 15^2 or 2a^2 = 225 or a^2 = 225/2 or a = 15sqrt(2)/2 and A = (1/2)(b)(h) = (1/2)(15sqrt(2)/2)(15sqrt(2)/2) = 225/4 square feet

OpenStudy (anonymous):